Accelerator Update for May 8, 2009 - May 11, 2009 |
ssThe Integrated Luminosity 5/4/09 to 5/11/09 was 59.05 pb-1. Friday May 8 The day shift began with Operations monitoring the Tevatron (TeV) store 7028 (210.06E30) and the Recycler stash (88.44E10), with the Antiproton Source stacking (11.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. Booster personnel worked on a Booster RF station (BRF2) and tuned for increased intensities despite decreasing RF voltages. Booster personnel repaired BRF2 and had it reconnected to the anode supply by 2 PM. Operations disabled the MTest beam at 6 PM. Operations terminated store 7028 at 10:48 PM. HEP shot setup followed and ran through to the end of the shift. Operations transferred antiprotons to the Recycler as appropriate. Saturday May 9 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (11.3E10) and the stack (43.1E10), and with NuMI and MiniBooNE taking beam. Shot setup was brought to a screaming halt by a D0 low beta quench. Recovery and investigation followed and continued through the rest of the midnight shift. During the quench recovery, TeV sectors A2 and C3 both showed mechanical dump switch failures. EE Support technicians were called in to investigate and repair the switch failures. LRF4 tripped on a driver RF low level fault and would not turn back on. Operators investigated and found that the driver regulator was missing its +5 volts. The Linac On-call tech came in to replace the bad power supply. MTest started taking beam shortly after 4 AM. Operations kept the TeV in a dry squeeze until noon and then made beam available for orbit checks. TeV personnel investigated the TeV kicker (C:PAK3) false pre/fail-to-fire trips during the dry squeeze. A MI power supply (M:HV102) developed an internal LCW leak that caused downtime at 10:45 AM. Near the flattop of the first orbit check, the TeV ramp dumped due to a TeV sector E1 lead failure. After consulting with TeV and cryo system personnel, the limits were adjusted. M:HV102 was repaired at 2:30 PM. The orbit checks were done by 3:30 PM. Operations had the TeV in shot setup by 4 PM. Operations established store 7035 by 5:45 PM with an average initial luminosity of 199.36E30. A Linac Klystron RF station (KRF3) tripped off several times requiring local resets. Operations disabled the MTest beam at 6:15 PM. Operations transferred antiprotons to the Recycler as appropriate. Sunday May 10 The midnight shift began with Operations monitoring store 7035 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. Operations established beam to MTest shortly after 4 AM. Operations terminated store 7035 at 9:15 AM. Operations put the TeV into HEP shot setup and established store 7036 at 11 AM with an average initial luminosity of 318.62E30. Operations disabled the MTest beam at 6:30 PM. There was some downtime due to KRF4 LCW flow trips. Operations transferred antiprotons to the Recycler as appropriate. Monday May 11 The midnight shift began with Operations monitoring store 7036 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. Operations terminated store 7036 at 12:02 AM. TeV sector B0 QPM communication problems were investigated following the store termination. Operations put the TeV through dry and wet squeezes. Operations put the TeV into shot setup at 2:43 AM. Operations established store 7039 at 4:07 AM with an average initial luminosity of 305.28E30. Operations established beam to MTest at 4 AM. Operations transferred antiprotons to the Recycler as appropriate. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 05/11/2009 email Fermilab |
| Security, Privacy, Legal | |